OpenStudy (warpedkitten):

If country A has 163,950 million more computers than country B and the total number of computers for both countries is 317,450 million, find the number of computers for each country.

OpenStudy (brainzonly):

Set up an algebraic equation: b+163,950=317,450

OpenStudy (warpedkitten):

So.. what I've tried to do... x = the amount of computers B has x + 163,950 = 317,450 So would 153,500 be the amount of computers for B?

OpenStudy (warpedkitten):

Didn't see you had posted that before me, but so my setup is correct?

OpenStudy (brainzonly):

Now... b+163,950-163,950=317,450-163,950 ^^Solve for B, which is what we're looking for b=153,500 Country B has 153,500 million computers, which would probably be 153500000000 computers! Yes, your setup is correct! You have a good day.

OpenStudy (warpedkitten):

Awesome. Thank you, and you too! :)

OpenStudy (brainzonly):

:D

OpenStudy (brainzonly):

Did I not answer your question?